OpenStudy (anonymous):

Verify the identity. cos(x+pi/2) = -sin x please help!

OpenStudy (anonymous):

There is an identity you can use, are you familiar with the co-function identities?

OpenStudy (anonymous):

Using the identities for sine you may arrive at proving the equation

OpenStudy (anonymous):

i think i have it hold on

OpenStudy (anonymous):

oh i got it @Johnbc but what do i do with the negative sign

OpenStudy (anonymous):

Well if you notice the cofunction of Sine would be this:\[Sin \theta=\cos(\frac{ \pi }{ 2 }- \theta)\]

OpenStudy (anonymous):

where our original shows that there is a plus in-between the angle and the pi/2 so we have to remove a multiple of -1 to get to: \[Sin \theta= - Cos (-\frac{ \pi }{ 2 } + \theta) \]

OpenStudy (anonymous):

so if you divide both sides by -1 and re-arrange whats inside the cosine you get

OpenStudy (anonymous):

\[-\sin \theta = \cos(\theta - \frac{ \pi }{ 2}) \] which now looks like your original function?

OpenStudy (anonymous):

cos(x+pi/2)= -sinx Use the cosine angle addition rule: cos(a+b) = cos(a)cos(b) - sin(a)sin(b) = cos(x)cos(pi/2)-sin(x)sin(pi/2) = cos(x)*0-sin(x)*1 = -sin(x)